One hundred identical mortgages are pooled together into a pass-through security. Each mortgage has a $100,000 principal, a fixed interest rate of 9% p.a. (paid monthly), and is fully amortized over a term of 30 years. Assume the bank charges servicing fees of 40 basis points and GNMA charges 10 bp to insure the timing of the payments. If investors expect no prepayment on these loans over the 30 years life of the pass-through, what should they pay for this package of mortgage backed securities (MBS), if they demand a return of 8%?
Answer #1

In the mortgage backed securities, all the returns are pass through. So the net return of pooled assets will be passed to the investor.

Pooled assets earn interest rate of 9%, and bank fees and GNMA charges cost 0.50%, that leaves us with net return of 8.5%.

Hence monthly net return = 100000 *8.5% /12 = 708.33

We can use the financial calculator to find the present value of the security:

N = 30*12 months = 360

I/Y = 8/12 = 0.67 (Per month)

PMT = 708.33

FV = 100000

Compute PV, we get 105203.96

Hence investor can pay 105203.96 in order to earn a return of 8%.